OpenStudy (anonymous):

The graph of f is perpendicular to the line whose equation is 9x-8y-24=0 and has the same y-int as this line Find slope intercept

OpenStudy (anonymous):

i got y=-9/8x-3

OpenStudy (anonymous):

to start out, but i don't know where to go from here

OpenStudy (anonymous):

oh wait yes i do

OpenStudy (anonymous):

oh wait no i dont. but i do know the slope is -9/8 and the y int is -3

OpenStudy (anonymous):

close, the slope is positive 9/8 but your y-int is correct \[9x - 8y - 24 = 0\] \[9x - 8y = 24\] \[-8y = -9x + 24\] \[y = \frac{ -9 }{ -8 }x + \frac { 24 }{ -8 }\] \[y = \frac{ 9 }{ 8 }x - 3\]

OpenStudy (anonymous):

Now it told us the y-int of f is the same as this line so we know so far that the graph of f is y = mx - 3 But we need a m and the line has to be perpendicular to the one given. Well a perpendicular line means its slope is the opposite reciprocal ( flip fraction and make it negative) of the one we know so the graph of f is y = -8/9x - 3

OpenStudy (anonymous):

oh i didn't flip it again, thats what i did wrong. thanks!